30. Suppose that in an economy, output per worker and capital per worker grew by 2.42%, and 2.48%, respectively. If the share of capital in output is 0.25:

(a) Calculate the total factor productivity.

(b) Is there a reason to believe that this economy is at steady state?


1
Expert's answer
2022-01-07T09:18:34-0500

(a)"A=\frac{Y}{K^aL^b}"

L=1

Y=1.0242

K=1.0248

a=0.25

b=0.75

"A=\frac{Y}{K^aL^b}=\frac{1.0242}{{1.0248^{0.25}}1^{0.75}}=1.017947" or 101.8%

(b)The economy is in a stage of slight growth, we can say that it is stable